# Sco X-1 as a continuous gravitational waves source: modelling the secular evolution using MESA

### v1

@@ 6.3.1 Sco X-1 detection probability in the case of crustal breakage
We estimate the probabilities $p_{r}$ and $p_{t}$ of detecting continuous gravitational waves from Sco X-1 crustal failure by respectively a radiometer search like (Abbott et al., 2021) and by a template-search like (Abbott et al., 2022).
@@ 7 Discussion
    Magnetic confinement of accreting matter
@@
accretion efficiency level and donor mass hugely impact gravitational wave emission: at $30\%$ accretion efficiency, we have no crustal breakage, therefore no gravitational wave emission. At higher accretion efficiencies, only a subset of Sco X-1 progenitors lead to the failure of the neutron star’s crust. In particular, for efficiencies between $40\%$ and $60\%$, we find a lower bound on the mass of the progenitor donor star leading the neutron star’s crust to break. This lower bound goes from $1.52\,M_{\odot}$ to $1.07\,M_{\odot}$, decreasing with increasing efficiency of the mass transfer.
For a few progenitor-accretion efficiency combinations the crust of the neutron star breaks while the system is already Sco X-1-like, so we can, in principle, observe the entire signal sweeping from $\approx 1000\,\textrm{Hz}$ to the torque-balance gravitational wave frequency. This high-frequency/loud gravitational wave phase cannot be detected by template-based searches, since the spin-down during this phase is too high. Radiometer searches or hidden Markov model searches (Melatos et al., 2021) are better suited to detect signals emitted during this phase. However, current detection probability is hindered by the fact that the detectable phase is very short-lasting and by the fact that the number of Sco X-1 configurations for which this happens is only a few. We find virtually no chances to detect such signals now.
Detection prospects with third-generation detectors are much brighter: not only is a much larger portion of the sweeping signal detectable, but also emission at torque balance, at least for the large majority of the possible signals. We estimate a probability of $41\%$ to detect gravitational waves from Sco X-1 by means of radiometer searches. If the leftover post-breakage ellipticity remains unchanged, the neutron star will keep emitting continuous gravitational waves in torque balance, that could be detected by the more sensitive template-based searches. We estimate the detection probability for these to be $45\%$.
@@ 7.1 Prospects
  • Spin-induced elastic deformations could be sourcing continuous gravitational waves from an accreting neutron star. Morales & Horowitz (2024) provide a model for this mechanism, with the resulting ellipticity depending on the degree of crustal anisotropy. The range of this quantity is not known. We have performed a preliminary investigation using the framework described in this paper, finding that the degree of crustal anisotropy necessary for a detection at current sensitivity must be higher than $1\%$. We will explore this mechanism in greater depth in the future.
  • The size of the total possible Sco X-1 parameter space affects the estimates of the detection probabilities. Whereas orbital period and mass of the donor star are observationally constrained, the range of possible neutron star deformations is not and it is solely based on estimates. Should more reliable estimates become available, they should be used to re-evaluate the detection prospects. We flag this as a point of caution in interpreting our detection probabilities.
  • We did not model the propeller effect, which could impact our results if at some point in its past evolution Sco X-1 did undergo a propeller phase. If that happened we would likely observe a transitional X-ray source, whereas Sco X-1 has been persistently accreting since it was first observed, suggesting that a propeller phase if at all, might be in Sco X-1’s future. Furthermore an appropriate modelling of the propeller effect will substantially increase the dimensionality of the parameter space, and might dominate probability estimates just due to sheer ignorance on its mechanisms.
@@ Acknowledgements
The inclusion of a Data Availability Statement is a requirement for articles published in MNRAS. Data Availability Statements provide a standardised format for readers to understand the availability of data underlying the research results described in the article. The statement may refer to original data generated in the course of the study or to third-party data analysed in the article. The statement should describe and provide means of access, where possible, by linking to the data or providing the required accession numbers for the relevant databases or DOIs.
@@ Appendix A Extra figures
Figure 7: Continuous gravitational wave amplitudes $h_{0}$ against their frequency (purple to yellow traces) stemming from compatible Sco X-1 systems in the case of magnetic confinement of accreted matter for different progenitors (indicated in the bottom left corner of each panel), accretion efficiencies, and ellipticities. To facilitate visualisation, we only display gravitational waves emitted at evolutionary stages where the system resembles Sco X-1. The chosen subset of binary parameters displayed is sufficient to appreciate the overall effect the various quantities have on the outcome of the simulations.Traces are color-coded according to the value of $b$, or equivalently, to the saturation ellipticity. Amplitude upper limits (dark grey curve) are taken from Abbott et al. (2022), assuming the spin axis of the neutron star to be orthogonal to Sco X-1’s orbital plane. The dash-dotted line is the $3G$ projected sensitivity of Abbott et al. (2022) (Equation 35 using $\mathcal{D}=62$ and $N=1$). The red dashed line is the nominal torque balance curve.

### v2

@@ 6.3.1 Sco X-1 detection probability in the case of crustal breakage
We estimate the probabilities $p_{r}$ and $p_{t}$ of detecting continuous gravitational waves from Sco X-1 crustal failure by respectively a radiometer search like (Abbott et al., 2021) and by a template-search like (Abbott et al., 2022b).
@@ 7 Discussion
    Magnetic confinement of accreted matter
@@
accretion efficiency level and donor mass hugely impact gravitational wave emission: at $30\%$ accretion efficiency, we have no crustal breakage, therefore no gravitational wave emission. At higher accretion efficiencies, only a subset of Sco X-1 progenitors lead to the failure of the neutron star’s crust. In particular, for efficiencies between $40\%$ and $60\%$, we find a lower bound on the mass of the progenitor donor star leading the neutron star’s crust to break. This lower bound goes from $1.55\,M_{\odot}$ to $1.07\,M_{\odot}$, decreasing with increasing efficiency of the mass transfer.
For a few progenitor-accretion efficiency combinations the crust of the neutron star breaks while the system is already Sco X-1-like, so we can, in principle, observe the entire signal sweeping from $\approx 1000\,\textrm{Hz}$ to the torque-balance gravitational wave frequency. This high-frequency/loud gravitational wave phase cannot be detected by template-based searches, since the spin-down during this phase is too high. Radiometer searches or hidden Markov model searches (Sun et al., 2018; Melatos et al., 2021) are better suited to detect signals emitted during this phase. However, current detection probability is hindered by the fact that the detectable phase is very short-lasting and by the fact that the number of Sco X-1 configurations for which this happens is only a few. We find virtually no chances to detect such signals now.
Detection prospects with third-generation detectors are much brighter: not only is a much larger portion of the sweeping signal detectable, but also emission at torque balance, at least for the large majority of the possible signals. We estimate a probability of $41\%$ to detect gravitational waves from Sco X-1 by means of radiometer searches. If the leftover post-breakage ellipticity remains unchanged, the neutron star will keep emitting continuous gravitational waves in torque balance, which could be detected by the more sensitive template-based searches. We estimate the detection probability for these to be $45\%$.
We note that these estimates rely on the fact that the crust breaks when $\Omega^{\textrm{br}}/2\pi\approx 550\,\textrm{Hz}$, which in turn depends on the assumed breaking strain value of $\sigma\approx 0.05$ (Morales & Horowitz, 2025). Crusts with a bigger breaking strain (up to $\sigma=0.1$ is possible in neutron stars Horowitz & Kadau (2009)) break at higher frequencies, which in our framework translates into fewer Sco X-1 compatible configurations that lead to crustal breakage, therefore negatively impacting the detection probability.
@@ 7.1 Prospects
  • Spin-induced elastic deformations could be sourcing continuous gravitational waves from an accreting neutron star. Morales & Horowitz (2024) provide a model for this mechanism, with the resulting ellipticity depending on the degree of crustal anisotropy. The range of this quantity is not known. We have performed a preliminary investigation using the framework described in this paper, finding that the degree of crustal anisotropy necessary for a detection at current sensitivity must be higher than $1\%$. We intend to explore this, and other promising mechanisms relevant to accreting neutron stars (e.g., thermal mountains, Li et al. (2025b, a)), in greater depth in the future.
  • The size of the total possible Sco X-1 parameter space affects the estimates of the detection probabilities. Whereas orbital period and mass of the donor star are observationally constrained, the range of possible neutron star deformations is not, and is based solely on estimates. Among the many assumptions made here, the range of ellipticities is what impacts most the detection probabilities. We have adopted a minimum value of $10^{-9}$ based on (Woan et al., 2018), but the smallest ellipticity could well be a few/several orders of magnitude smaller (Mastrano et al., 2011), leading to detection probabilities correspondingly smaller. For example, if the smallest ellipticity happens to be $\varepsilon_{\textrm{min}}\approx 10^{-12}$, which is sometimes used as a benchmark for the lowest deformation in magnetically deformed neutron stars (Haskell et al., 2008), the parameter space available to Sco X-1 will be $\approx 1.7$ times bigger, and detection probabilities will decrease roughly of the same factor given that ellipticities smaller than $10^{-9}$ are essentially beyond detectability. We flag this as a point of caution in using our detection probabilities.
  • We did not model the propeller effect, which could impact our results if, at some point in its past evolution, Sco X-1 did undergo a propeller phase. If that happened, we would likely observe a transitional X-ray source, whereas Sco X-1 has been persistently accreting since it was first observed, suggesting that a propeller phase, if at all, might be in Sco X-1’s future. Furthermore, an appropriate modelling of the propeller effect will substantially increase the dimensionality of the parameter space, and might dominate probability estimates just due to sheer ignorance of its mechanisms.
    On the other hand, it is worth reflecting on the more realistic scenario of a time-varying accretion efficiency. At the onset of mass accretion we expect $\eta\approx 100\%$, with $\eta$ later on decreasing due to the many non-conservative mass transfer processes that can take place in Sco X-1, as mentioned in Section 3.1. It might be that the set of $P-M^{d}$ configurations evolving into Sco X-1 in this more complex scenario, would not be too far off from those found in the simpler case of constant $\eta$, as long as the time-averaged accretion efficiency is not $\lesssim 30\%$. We say this because we find that the sets of progenitors obtained at different fixed accretion efficiencies between $40\%\leq\eta\leq 80\%$, are roughly similar to each other. Time-dependent accretion efficiency might instead significantly affect which Sco X-1 progenitors lead to the failure of the neutron star’s crust. If indeed the $\eta\approx 100\%$ phase lasts long enough, crustal failure may also occur for those progenitor systems for which we find the crust does not break in the simplified case of constant $\eta$.
@@ Acknowledgements
MESA inlists and data required to calculate detection probabilities are available on the official website of the Continuous Gravitational Waves group at the MPI/AEI Hannover (Pagliaro et al., 2025a), as well as on the journal’s webpage. MESA inlists will also be available on ZENODO.
@@ Appendix A Extra figures
Figure 7: Continuous gravitational wave amplitudes $h_{0}$ against their frequency (purple to yellow traces) stemming from compatible Sco X-1 systems in the case of magnetic confinement of accreted matter for different progenitors (indicated in the bottom left corner of each panel), accretion efficiencies, and ellipticities. To facilitate visualisation, we only display gravitational waves emitted at evolutionary stages where the system resembles Sco X-1. The chosen subset of binary parameters displayed is sufficient to appreciate the overall effect the various quantities have on the outcome of the simulations.Traces are color-coded according to the value of $b$, or equivalently, to the saturation ellipticity. Amplitude upper limits (dark grey curve) are taken from Abbott et al. (2022b), assuming the spin axis of the neutron star to be orthogonal to Sco X-1’s orbital plane. The dash-dotted line is the $3G$ projected sensitivity of Abbott et al. (2022b) (Equation 35 using $\mathcal{D}=62$ and $N=1$). The red dashed line is the nominal torque balance curve.